Please help me solve:
John is 4 years older than his sister, and 6 years ago he was twice his sisters age then. How old is each now?

Answer by oberobic(2304)

J = John's current age
S = John's sister's current age
.
John is 4 years older than his sister
J = S+4
.
6 years ago their ages were:
J-6
S-6
.
6 years ago John was twice as old as his sister
J-6 = 2(S-6)
J = 2S -12 +6
J = 2S -6
.
substitute
.
J = S+4
2S -6 = S+4
S = 10
.
J = S+4
J = 14
.
Check the solution to be sure it is the answer.
.
Six years ago...
J-6 = 8
S-6 = 4
8 = 2*4
Correct.
.
Answer: John is 14 years old, and his sister is 10.
.
Done.